OpenStudy (anonymous):

solve for x:2x + 6=4x +20

OpenStudy (doc.brown):

\[2x+6=4x+20\]\[2x+6-6=4x+20-6\]\[2x=4x+20-6\]\[2x-4x=4x-4x+20-6\]\[2x-4x=20-6\]\[-2x=14\]\[\frac{-2x}{-2}=\frac{14}{-2}\]\[1x=-7\]\[x=-7\]check\[2x+6=4x+20\]\[2(-7)+6=4(-7)+20\]\[-14+6=-28+20\]\[-8=-8\] yup